BE1-5 Use the accounting equation to answer these independent questions: (a) The shareholders equity of Sansom Corporation
Answer #1

a)

Stockholders' equity = $120,000

Total liabilities = $55,000

Total assets = ?

Total assets = Total liabilities + Stockholders' equity

= 55,000 + 120,000

= $175,000

b)

Total liabilities = $170,000

Share capital = $100,000

Retained earnings = $90,000

Stockholders' equity = Share capital + Retained earnings

= 100,000 + 90,000

= $190,000

Total assets = Total liabilities + Stockholders' equity

= 170,000 + 190,000

= $360,000

c)

Total assets = $150,000

Share capital = $50,000

Retained earnings = $25,000

Stockholders' equity = Share capital + Retained earnings

= 50,000 + 25,000

= $75,000

Total assets = Total liabilities + Stockholders' equity

150,000 =  Total liabilities + 75,000

Total liabilities = $75,000

d)

Total assets = $500,000

Total liabilities = $250,000

Total assets = Total liabilities + Stockholders' equity

500,000 = 250,000 + Stockholders' equity

Stockholders' equity = $250,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.